OpenStudy (anonymous):

Find the amount of interest and the monthly payment for the loan. Purchase a car for $42,600 at 2.7% add-on rate for 5 years. (Round your answers to the nearest cent.)

OpenStudy (anonymous):

Apply the formula \[ \frac{P r}{12 \left(1-\left(\frac{r}{12}+ 1\right)^{-m}\right)} \] Where P=42600 r=.027 m=60 You cam finish it now.

OpenStudy (anonymous):

You can also do it on my website http://saab.org on the last link